OpenStudy (anonymous):

Find the derivative of f(x) = 5 divided by x at x = -1.

OpenStudy (nincompoop):

use quotient rule then substitute the value at x=-1

OpenStudy (anonymous):

I dont understand at all

OpenStudy (kirbykirby):

\[ \frac{5}{x}=5x^{-1}\] and use the power rule, it's easier than using the quotient rule.

OpenStudy (anonymous):

is it 5?

OpenStudy (kirbykirby):

There's something missing

OpenStudy (kirbykirby):

\[ \frac{d}{dx}5x^{-1}=5(-1)x^{-2}\], now evaluate at \(x=-1\)

OpenStudy (nincompoop):

nothing is easier than anything it is a matter of technique

OpenStudy (kirbykirby):

Well I meant easier in the sense that you perform less steps, and you don't have to worry about remembering an additional rule if you don't want. (All quotients can technically be differentiated with the product + chain rule)

OpenStudy (nincompoop):

haaaaaa?

OpenStudy (kirbykirby):

For example: \[ \frac{4x+1}{2x^2+2}=(4x+1)(2x^2+2)^{-1}\]

OpenStudy (nincompoop):

laughing out loud I am messing with you. Again, it is a matter of technique and whatever the person is proficient at doing

OpenStudy (kirbykirby):

^true

OpenStudy (nincompoop):

well you've just set yourself up, young man I can do that fast in quotient rule than you can in product

OpenStudy (kirbykirby):

I suppose, but again, whatever technique suits the person. And I think the power rule above is faster than using the quotient rule :P